potassium has a bdd structure with nearest neighbour distance of 4.54 arngstrong. its density will be in kgm-3.
options:a)4.54  b)808  c)852  d)910 .
  
 
 
 

Grade:12

1 Answers

Arun
6 years ago
For bcc
d = \sqrt3*  a/2
a = 2d /\sqrt3
a = 2*4.54/1.732
= 522 Pico metre
 
Now density = z *M/N*a³
= 2*39/6.023 *10²³ *(522)³ * 10-30
Density = 910 kg /m³